OpenStudy (anonymous):

The expression below is the factorization of what trinomial? -1(x-6)(x+8)

OpenStudy (anonymous):

x=6 x=-8

OpenStudy (anonymous):

(x+a)(x+b) = x^2+ax+bx+ab

OpenStudy (anonymous):

= x^2+(a+b)x+ab

OpenStudy (anonymous):

I assume you know how to multiply polynomials. when I multiplied (x+a) with (x+b), I got a polynomial with three terms, otherwise known as a trinomial.